zarkam21 (zarkam21):

Solve the system of equations. Enter your answer as an ordered pair. 5x+3y=31 2x+3y=25

Mehek (mehek14):

first eliminate y by subtracting both equations

Mehek (mehek14):

5x + 3y = 31 - 2x + 3y = 25 --------------- 3x + 0 = 6 x = 2 now substitute 2 in the equations

zarkam21 (zarkam21):

(2,7) ??

Mehek (mehek14):

yup